- Press Release
- Dec 7, 2022
NASA Space Station On-Orbit Status 22 January 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below. Saturday – first weekend off-duty day for the crew (except for standard maintenance and EVA preparation tasks). Day 101 of Expedition 10.
CDR/SO Chiao and FE Sharipov performed the regular weekly 3-hr. task of thorough station cleaning. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]
In more software upgrading, Sharipov first configured the Russian Laptop 2 (TP2) for upgrading the firmware of BSV-M2, the second time synchronization unit (of two) on its system power panels (PPS). After consulting with the ground, he completed the firmware step-up for BSV-M2. [BSV-M1 was upgraded on 1/10.]
As further preparation for the EVA-12 on 1/26, the crew set up and reconfigured the spherical Transfer Compartment (PkhO) of the Service Module (SM) for isolation and EVA activities, including disassembly of the air vent duct to the DC1. [The hatch from PkhO to the DC1 docking module vestibule will be closed during the spacewalk while the second hatch on DC1-side remains open. The hatch to the FGB tunnel will stay open, but the second hatch between vestibule and FGB will be closed.]
In the US Airlock, Leroy initiated the recharging of the four EVA helmet light batteries (#1009, #1016, #1019, #1020) in the BSA (Battery Stowage Assembly). [Chiao will terminate the charging tomorrow and install the batteries into the helmet light assemblies on the Orlan-Ms. At that time, he will also test the function of the flood and spotlights in the helmet light assemblies.]
Salizhan did the routine maintenance of the SM’s SOZh environment control & life support system, including ASU toilet system service and replacements.
Leroy rebooted the SSC OCA (station support computer/orbital communications adapter) router laptop, a periodic requirement.
The CDR also collected the periodic reading of the cabin air’s current CO2 partial pressure in the SM and Lab, using the U.S. CDMK (carbon dioxide monitoring kit), for calldown (along with the battery status) for use in trending analyses.
The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser and VELO cycle with bungee cord load trainer. Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 4 of a new set) and one hour on VELO.
Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
At ~5:30am EST, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.
Last night, the ground completed the 4B1 battery capacity test. Battery set 4B1 was taken through a complete discharge cycle and then recharged. The BCDU-1 (Battery Charge/Discharge Unit #1) is back online and Channel 4B is once again operating at full capability.
Overnight from 1/20 to 1/21, the Elektron O2 generator’s primary and backup pumps failed. The suspected cause was bubbles in the system. The crew performed troubleshooting by purging the line of air bubbles. The Elektron was restarted and is now operating nominally in 32-amp mode.
A new task has been added to the EVA. The crew will perform inspection and photography of the Elektron, Vozdukh and Micropurification Unit (BMP) gas vent deflectors on the SM’s exterior.
Weekly Science Update (Expedition Ten — 12th):
GASMAP: Nothing new.
Human Research Facility/Workstation (HRF WS): Continuing.
Advanced Ultrasound (ADUM): Nothing new.
Hand Posture Analyzer (HPA: Nothing new.
In-Space Soldering Investigation (ISSI): Operations are complete.
Pore Formation and Mobility Investigation (PFMI): Nothing new.
Space Acceleration Measurement System (SAMS): Nothing new.
Microgravity Acceleration Measurement System (MAMS): Nothing new.
Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.
Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.
Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.
Binary Colloidal Alloy Test-3 (BCAT-3): Nothing new.
Renal Stone (RS): Nothing new.
Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.
Foot/Ground Reaction Forces during Space Flight (FOOT): Nothing new.
Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.
Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): Nothing new.
Biotechnology Specimen Temperature Controller (BSTC): Nothing new.
Yeast Group Activation Packs (Yeast GAP): Nothing new.
Earth Knowledge Acquired by Middle School Students (EarthKAM): Nothing new.
Earth Science Toward Exploration Research (ESTER): Nothing new.
Serial Network Flow Monitor (SNFM): Nothing new.
Fluid Merging Viscosity Measurement (FMVM): Nothing new.
Space Experiment Module (SEM): Nothing new.
Viscous Liquid Foam–Bulk Metallic Glass (Foam): Nothing new.
Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): Nothing new.
Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.
Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.
Miscible Fluids in Microgravity (MFMG): Nothing new.
Educational Payload Operations (EPO): Nothing new.
Capillary Flow Experiment (CFE): Nothing new.
Crew Earth Observations (CEO): The ground has received, and at least preliminarily reviewed, 7,482 of Increment 10 CEO images to date. A pair of recent images of the northwestern coast of Sumatra illustrating dramatic tsunami damage will be published on NASA’s Earth Observatory website this weekend. The image pair, taken during the same pass, also demonstrated the use of sun glint in Astronaut photography to detect standing water in comparable views. Thanks to the crew for their fine technique and patience in acquiring these images. The ground also spotted “some of the finest long-lens images of downtown Houston we’ve seen in some time. The roof of Minutemaid Park was clearly open! Wow!”
Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Calcutta, India (this nadir pass provided an opportunity for mapping of this megacity. Mapping swaths from the urban-rural fringe across the urban core are particularly useful for quantifying urban growth and spatial landscape structure), Lima, Peru (weather was predicted to be clear along the coastline for imagery of this rarely cloud-free city. Mapping swaths along the eastern urban-rural fringe [close to the Andes] are useful for assessment of geohazards and ecological disturbance resulting from land use change), Internal waves, New Zealand (this overpass provided an opportunity for internal wave photography. Looking to the right of track to the northwest of North Island for the sunglint point), and Jarvis Island, equatorial Pacific (weather was predicted to be clear for high-resolution reef photography. Images of the reef are useful for assessment of its extent and ecological health when compared with previous imagery).
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- EVA-12 — 1/26/05 (hatch opening 2:27am EST)
- Progress 16P undocking & destructive reentry — 2/27/05;
- Progress 17P launch — 2/28/05.
- EVA-13 — 3/25/05;
- Soyuz 10 S launch — 4/15/05;
- Soyuz 9S undock — 4/25/05 (after 193 days on orbit, 191 days on board ISS).
ISS Orbit (as of this morning, 7:20am EST [= epoch]):
- Mean altitude — 358.4 km
- Apogee height — 364.9 km
- Perigee height — 351.9 km
- Period — 91.71 min.
- Inclination (to Equator) — 51.65 deg
- Eccentricity — 0.0002703
- Solar Beta Angle — -26.6 deg (magnitude decreasing)
- Orbits per 24-hr. day — 15.70
- Mean altitude loss in last 24 hours — 100 m
- Revolutions since FGB/Zarya launch (Nov. 98) — 35278
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.